OpenStudy (anonymous):

How do I graph the polynomial function of f(x)=x^4-3?

OpenStudy (anonymous):

f(x) = y y = x^4 - 3 you can graph the plot by substituting values in x For x = 0, y = (0)^4 -3 = -3 For x = 1, y = (1)^4 - 3 = -2 For x = -1, y = (-1)^4 - 3 = -2 and so on

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

Anytime :)

OpenStudy (anonymous):

How about if I am using a polynomial like this: f(x)=x3-4x+1?